OpenStudy (anonymous):

derivative: 1. g(t)=√√t+1 +1

OpenStudy (amistre64):

you need to better clarify the function

OpenStudy (anonymous):

\[g(t) = \sqrt{\sqrt{t+1}}+1\]

OpenStudy (anonymous):

the +1 is included inside the big sqrt :D

OpenStudy (amistre64):

\[g(t) = \sqrt{\sqrt{t+1}+1}~~\]

OpenStudy (anonymous):

yep thats it

OpenStudy (amistre64):

its best to work with powers, sqrt = ^1/2 other than that, it looks like power and chain rules to apply

OpenStudy (amistre64):

have you covered the derivative rules yet, or is this one of thos first principles thing with the limit of a slope?

OpenStudy (anonymous):

after ^1/2 apply chain rule?

OpenStudy (amistre64):

\[((t+1)^{1/2}+1)^{1/2}\] we have a function in a function, so chain rule applies:\[f(g)=f'(g)~g'\]

OpenStudy (amistre64):

let f = a^(1/2) f' = a'/[2a^(1/2)] but a = (t+1)^(1/2) + 1, so a' = 1/[2(t+1)^(1/2)] agreed?

OpenStudy (anonymous):

\[1/2(\sqrt{t+1}+1)x((t+1)^{1/2}) \]

OpenStudy (amistre64):

\[((t+1)^{1/2}+1)^{1/2}\] \[\frac12((t+1)^{1/2}+1)^{-1/2}~*~((t+1)^{1/2}+1)'\] \[\frac12((t+1)^{1/2}+1)^{-1/2}~*~\frac12(t+1)^{-1/2}\] \[\frac14((t+1)^{1/2}+1)^{-1/2}(t+1)^{-1/2}\] \[\frac14[(t+1)^{1/2}+1)(t+1)]^{-1/2}\]

OpenStudy (amistre64):

or written in sqrts:\[\frac{1}{4\sqrt{(\sqrt{t+1}+1)(t+1)}}\]

OpenStudy (anonymous):

can you help me with another one?